If x=3 and x=-5, then form a quadeatic equation. Can somebody help me?

Found 2 solutions by checkley71, 067isa:
Answer by checkley71(8403)

X=3 OR X-3=0
X=-5 OR X+5=0
(X-3)(X+5)
X^2+2X-15 IS THE QUADRATIC EQUATION WITH SOLUTIONS FOR X=3 & X=-5

Answer by 067isa(13)

x-3=0 and x+5=0
by multiplying
(x-3)(x+5)=0
X^2+5X-3X-15=0
collect the like terms
X^2+2X-15=0